OpenStudy (mathmath333):

Doctors advised Renu, a chocolate freak, not to take more than 20 chocolates in one day. when she went out to the market to buy her daily quota, she found that if she buys chocolates from the market complex she would have to pay rs.3 more for the same number of chocolates than she would have spent had she bought them from her uncle Scrooge's shop, getting two sweets less per rupee. she finally decided to get them from her uncle's shop paying in only one rupee coins 1. how much would she have spent at the market complex? a. 6 b 12 c 9 d 5

OpenStudy (mathmath333):

@amistre64

OpenStudy (amistre64):

how are you parsing the information so far?

OpenStudy (amistre64):

seems like a lot of words to confuse people with to me

OpenStudy (mathmath333):

yes its confusing

OpenStudy (amistre64):

i cant make much sense of it at the moment

OpenStudy (mathmath333):

i have found that the number of chocolates is a multiple of 6

OpenStudy (amistre64):

did she pay 1 rupee for 20 sweets in the end?

OpenStudy (mathmath333):

no it means like she gave money in one cent coins only

OpenStudy (amistre64):

if she buys X chocolates from the market complex she would have to pay 3r more for the same X number of chocolates than she would have spent had she bought [X number of chocolates] from her uncle Scrooge's shop, getting two sweets less per rupee. M$/X = (U$ + 3)/X and im thinking of some ratio of X:X-2

OpenStudy (amistre64):

just trying to iron out a thought (X-2)/(U+3) = X/U UX - 2U = UX+3X - 2U = 3X U = -3/2 X M - 3 = -3/2 X M = -3/2 X + 3 M = 3/2 X if X is a multiple of 6 M = 9n none of this really makes a lick of sense to me but its what im coming up with. im going to say we need a multiple of 9 but i have no way to verify that any of this is correct lol

OpenStudy (amistre64):

6 does not divide 20 evenly tho

OpenStudy (amistre64):

1 20 2 10 4 5

OpenStudy (amistre64):

yeah, im going with i havent a clue